OpenStudy (anonymous):

x^1/2 times x^1/2 times x^1/2 .

OpenStudy (psymon):

When multiplying, if all the numbers you're multiplying have the same base, you simply add the exponents together. All of these 3 have a base of x, so you can multiply them by just adding together the 3 exponents and making that as your new power.

OpenStudy (zzr0ck3r):

\[a^b*a^c=a^{b+c}\]

OpenStudy (anonymous):

\[x^{\frac{1}{2}} \times x^{\frac{1}{2}} \times x^{\frac{1}{2}} = x^{\frac{1}{2}+\frac{1}{2}+\frac{1}{2}}\] \[= x^{\frac{3}{2}}\]